# A Survey of Model Architectures in Information Retrieval

### v2

@@ 7 Large Language Models for IR
In conclusion, the adoption of LLMs in IR represents more than a simple increase in model scale. While they certainly serve as more powerful feature extractors within existing bi-encoder and cross-encoder frameworks, their unique generative and instruction-following abilities are forging entirely new architectural paradigms like generative retrieval and zero-shot listwise reranking. As these models continue to evolve, significant challenges and open questions regarding their efficiency, scalability, and trustworthiness remain, paving the way for future research (§˜8).
@@ 8 Emerging Directions and Challenges
IR systems have become crucial across diverse domains, from retrieval-augmented language modeling (Khandelwal et al., 2020; Borgeaud et al., 2022) to applications in agents (Wu et al., 2023a; Wang et al., 2024a), code generation (Wang et al., 2024c; Zhang et al., 2023a), robotics (Anwar et al., 2024), medicine (Jeong et al., 2024), and protein research (Jumper et al., 2021), inter alia. These developments present new challenges for IR research. Drawing from the evolution of IR architectures (§§˜3, 4, 5, 6 and 7), we examine emerging trends, open problems, and potential research directions. We structure our discussion around three key areas: advancing the core components of IR models, adapting to the new paradigm of retrieval for AI, and tackling the pragmatic challenges of real-world deployment.
@@ 8.1 Advancing the Core Components of IR Models
  • Data and training efficiency. Current transformer-based IR models demand extensive training data (Fang et al., 2024), making them impractical for many real-world applications. Developing architectures that can learn effectively from limited data or in a few-shot/zero-shot setting remains crucial. Additionally, models should support parallel processing and low-precision training to reduce costs and accelerate convergence (Nvidia, 2021; Fishman et al., 2024; Liu et al., 2024a).
  • Inference optimization. Real-time applications like conversational search (Mo et al., 2024b) and agent-based systems (Yao et al., 2023) require efficient handling of variable-length queries, necessitating advanced compression and optimization techniques for both retriever backbones and index structures (Dettmers & Zettlemoyer, 2023; Kumar et al., 2024; Warner et al., 2024; Bruch et al., 2024; Xu et al., 2025a, inter alia).
  • Multimodality and multilinguality. Future IR systems must handle diverse content types including images (Ma et al., 2024b), audio (Pusateri et al., 2024), and structured data (Tan et al., 2024; Edge et al., 2024) as well as multilinguality beyond English (Zhang et al., 2023c; Enevoldsen et al., 2025). Recent advances in multimodal, multilingual retrieval (Ma et al., 2024b; Wei et al., 2025; Yu et al., 2024b; Huang et al., 2024b; Jiang et al., 2025, inter alia) and structured data processing (Li et al., 2023c; 2024) have demonstrated promises.
  • Transformer alternatives. While transformers have dominated recent IR research, their quadratic complexity in attention computation remains a significant bottleneck. Recent advances in linear RNNs (Peng et al., 2023; 2024; Qin et al., 2024c), state space models (Gu & Dao, 2024; Dao & Gu, 2024), and linear attention (Katharopoulos et al., 2020; Yang et al., 2024) offer alternatives with theoretical linear complexity. Although preliminary studies (Xu, 2024; Zhang et al., 2024; Xu et al., 2025c; b) show limited gains compared to optimized transformers, developing efficient alternatives architectures for transformers could revolutionize large-scale information retrieval.
Ultimately, strong foundation models have proven crucial for IR performance (Neelakantan et al., 2022; Ma et al., 2024c, inter alia). As IR applications expand, developing foundation models that balance computational efficiency with robust performance across tasks and modalities emerges as a key research priority.
@@
As discussed in §˜6, cross-encoders provide complex non-linear relevance estimation but are computationally expensive. In contrast, bi-encoder architectures used in dense and sparse retrieval rely on linear similarity functions (e.g., inner product) to enable fast retrieval through nearest neighbor search and inverted indexing. Balancing complex relevance matching and scalable retrieval remains challenging. ColBERT (Khattab & Zaharia, 2020) addresses this by using document representation matrices with MaxSim operations, while recent work (Killingback et al., 2025) explores Hypernetworks (Ha et al., 2022) to generate query-specific neural networks for relevance estimation. The design of flexible yet scalable relevance estimators remains an active research direction.

### v3

@@ 7.4 Broader Ecosystem and Concluding Remarks
In conclusion, the adoption of LLMs in IR represents more than a simple increase in model scale. While they certainly serve as more powerful feature extractors within existing bi-encoder and cross-encoder frameworks, their unique generative and instruction-following abilities are forging entirely new architectural paradigms like generative retrieval and zero-shot listwise reranking. However, the advancement of IR architecture is not driven solely by the pursuit of superior semantic matching capabilities. The practical deployment of these systems — ranging from lightweight encoders to massive LLMs — necessitates architectures that can withstand rigorous efficiency constraints, handle diverse data modalities, and ensure reliability. We examine these cross-cutting architectural adaptations in the following section.
8 Architectures for Diverse Scenarios and Constraints
The evolution of IR models described in previous sections — from vector space models to LLMs — primarily traces the pursuit of better semantic matching for English text. However, deploying these models in real-world environments requires navigating complex scenarios and constraints beyond pure textual relevance. These include handling diverse data modalities and languages, balancing the inherent tradeoff between accuracy and latency, and ensuring model reliability through calibration. In this section, we review how IR architectures are adapted to meet these specific requirements. Across these settings, architectural choices such as representation granularity and modularity, serve as the primary mechanisms to balance task-specific constraints with scalable retrieval.
8.1 Architectures for Multimodal and Multilingual Data
8.1.1 Multilingual and Cross-Lingual Architectures
Problem Definitions and Retrieval Settings.
Although often used interchangeably, Multilingual Information Retrieval (MLIR) and Cross-Lingual Information Retrieval (CLIR) correspond to distinct retrieval scenarios. CLIR refers to the setting in which a user issues a query in a source language and retrieves documents written in a different target language (Oard & Dorr, 1998; Fluhr et al., 1999). This distinction has direct architectural implications: CLIR requires explicit cross-language alignment at query time, whereas MLIR emphasizes building shared or interoperable representations during indexing. The central architectural challenge in CLIR is bridging the lexical and semantic gap between the query and document language spaces (Goworek et al., 2025). In contrast, MLIR is a broader paradigm in which a system indexes and searches over document collections containing multiple languages, potentially serving queries in any of them (Oard & Dorr, 1996; Oard et al., 1999; Fluhr et al., 1999). Architecturally, CLIR can be viewed as a special case of MLIR that explicitly requires cross-language alignment at retrieval time.
Early Translation-Centric Architectures.
The architectural foundations of cross-lingual retrieval date back to early work on the Vector Space Model (Salton et al., 1975), where cross-language retrieval was framed as a synonymy problem. Early systems relied on bilingual thesauri to map query terms into a shared conceptual space prior to retrieval, treating translation as a distinct pre-processing step (Salton, 1969; 1970). Architecturally, these systems isolated linguistic complexity into a separate translation component, leaving the retrieval engine itself unchanged and monolingual. When high-quality thesauri were available, such architectures could approach monolingual retrieval performance; however, they were brittle due to limited vocabulary coverage, poor handling of polysemy, and an inability to model multi-word expressions or contextual meaning.
Statistical and Representation-Based Cross-Lingual Models.
The availability of large parallel corpora in the 1990s, such as the Canadian Hansards [^5] https://en.wikipedia.org/wiki/Hansard, enabled a shift from static dictionaries to statistically grounded architectures. A major departure from direct translation was introduced by Cross-Language Latent Semantic Indexing (CL-LSI), which projected documents and queries from different languages into a shared, language-independent latent space using Singular Value Decomposition over parallel data (Dumais et al., 1997). In parallel, Statistical Machine Translation (SMT) became a dominant architectural component in CLIR systems, with retrieval pipelines adopting either query translation or document translation using probabilistic alignment models such as the IBM Models (Brown et al., 1993). Both CL-LSI and SMT-based pipelines preserved monolingual retrieval backends, differing primarily in whether cross-lingual alignment was achieved through latent semantic projection or probabilistic translation. These systems established the translation-based retrieval paradigm, where the retrieval engine itself remained monolingual and linguistic complexity was isolated within the translation module.
Multilingual Transformers and End-to-End Retrieval.
The introduction of multilingual pre-trained transformers, including mBERT and XLM-R, marked a fundamental architectural shift away from explicit translation toward shared semantic representation learning (Devlin et al., 2019; Conneau et al., 2020a; MacAvaney et al., 2019a). Trained on large-scale multilingual corpora, these models enabled a single encoder to represent queries and documents across languages, substantially reducing dependence on parallel data (Conneau et al., 2020b; Feng et al., 2022; Shi et al., 2020; Goswami et al., 2021). However, general-purpose multilingual encoders often underperform in retrieval settings due to insufficient cross-lingual alignment induced during pretraining (Zhang et al., 2022; Elmahdy et al., 2024), which has been the focus of subsequent research works.
Per-Language Modules.
To improve the performance of multilingual pre-trained transformers on low-resource languages, a line of works proposed to add language-specific adapters to enable zero-shot or few-shot cross-lingual transfer. Early explorations in NLP community focus on classical NLP tasks such as dependency parsing and named entity recognition (Üstün et al., 2020; Pfeiffer et al., 2020; Artetxe et al., 2020; Pfeiffer et al., 2021). In the context of IR, Litschko et al. (2022) compared the performance of adapter-based approaches and Sparse-Fine-Tuning Masks (Ansell et al., 2022) to NMT-based approaches and reported their efficacy in both performance and training efficiency.
Alignment-Focused and Modern Architectures.
Tackling the same challenge of per-language modules approaches, specialized multilingual retrieval models such as LaREQA, InfoXLM, and LaBSE enforce tighter alignment between semantically equivalent cross-lingual pairs using parallel corpora and contrastive objectives (Roy et al., 2020; Chi et al., 2021; Feng et al., 2022; Huang et al., 2024b). Notably, these approaches typically preserve the standard transformer backbone, focusing architectural innovation on representation learning objectives rather than structural redesign. Recent work further refines end-to-end multilingual retrieval through improved data curation, supervision, and knowledge distillation, continuing the shift away from multi-stage translation pipelines toward unified models that directly match meaning across languages (Zhang et al., 2023d; Yang et al., 2024).
8.1.2 Multimodal Architectures
From Shallow Fusion to Deep Joint Embeddings.
Early multimodal retrieval architectures relied on modality-specific feature extraction followed by shallow fusion, with text typically represented using bag-of-words models and images encoded via hand-crafted descriptors such as scale-invariant feature transform (SIFT) (Lowe, 1999). To enable cross-modal comparison, projection-based methods such as canonical correlation analysis (CCA) (Hotelling, 1992) mapped heterogeneous features into a shared embedding space, establishing the foundational principle that multimodal retrieval requires learning semantic correspondences across modalities rather than treating them independently (Rasiwasia et al., 2010; Sharma et al., 2012; Gong et al., 2014; Ranjan et al., 2015). The transition to deep learning replaced manual feature engineering with end-to-end representation learning, exemplified by early visual-semantic embedding models such as DeViSE (Frome et al., 2013; Wang et al., 2016). Sentence encoders incorporating syntactic structure, including Dependency Tree Recursive Neural Networks (DT-RNNs), further improved alignment by modeling relational semantics (Socher et al., 2014). Fragment-level architectures extended this paradigm by decomposing images and sentences into finer-grained units and aligning them with structured max-margin objectives, enabling both global and local cross-modal reasoning (Karpathy et al., 2014). Adversarial frameworks such as ACMR subsequently refined joint embeddings by introducing nonlinear projections and modality-invariant regularization (Wang et al., 2017). We refer to Wang et al. (2024b) for a comprehensive survey.
Dual-Encoder Contrastive Models and Controlled Interaction.
Large-scale vision–language datasets enabled a paradigm shift toward bi-encoder architectures trained with contrastive objectives. CLIP and ALIGN independently encoded images and text and aligned them through contrastive learning on hundreds of millions to billions of image–text pairs, establishing scalable and efficient retrieval backbones (Radford et al., 2021; Jia et al., 2021; Wei et al., 2025). This architectural separation facilitated efficient indexing and retrieval and rapidly generalized to additional modalities including video, audio, depth, and sensor data (Girdhar et al., 2023; Chen et al., 2023b; Kong et al., 2025). Subsequent refinements explored the efficiency–expressivity tradeoff within this paradigm, most notably through late-interaction architectures such as ColBERT, which replaced single-vector embeddings with multi-vector representations to enable token-level matching while preserving much of the efficiency of bi-encoders (Khattab & Zaharia, 2020; Faysse et al., 2025; Wan et al., 2025). Together, these models established a dominant architectural family centered on modality-separated encoding with limited but scalable interaction.
Rich Fusion, Hierarchical Interaction, and Unified Architectures.
Beyond bi-encoders, multimodal retrieval architectures increasingly incorporated sophisticated mechanisms to support complex reasoning, particularly in video–text retrieval (Chen et al., 2020a). Transformer-based and hierarchical models decomposed alignment into global-to-local or multi-granular stages, combining cross-modal attention with temporal and semantic structure (Gabeur et al., 2020; Liu et al., 2021; Zhang et al., 2023b; Gorti et al., 2022). Hybrid designs integrated CLIP-style encoders with cross-modal fusion or temporal alignment modules, balancing pretrained representations with task-specific interaction (Portillo-Quintero et al., 2021; Fang et al., 2021). These architectures substantially increase computational cost, often restricting their use to reranking or small candidate sets.
Recent architectures further differentiate between multi-encoder designs that preserve modality-specific encoders and single-encoder models employing full cross-attention at higher computational cost (Li et al., 2019; 2022b; Zhai et al., 2023; Kim et al., 2025). At the same time, multimodal large language models (MLLMs) and multi-agent retrieval frameworks unify retrieval, reasoning, and generation within a single system, while modality-preserving and any-to-any architectures emphasize flexible interaction without collapsing modality structure (Liu et al., 2023; Xie et al., 2024a; Liu et al., 2025b; Xu et al., 2025a; Ju et al., 2025).
Remarks.
From a structural perspective, multimodal retrieval architectures can be broadly grouped into: (1) shallow fusion and projection-based models, (2) deep joint-embedding architectures with global or fragment-level alignment, (3) contrastively trained dual-encoder models emphasizing scalability, (4) late-interaction hybrids balancing efficiency and expressivity, and (5) cross-attention and MLLM-centric architectures enabling deep fusion and unified multimodal reasoning. This progression reflects a recurring architectural tension between interaction richness and computational efficiency.
8.2 Performance-Efficiency Tradeoffs and the Role of Architectural Choices
Retrieval systems face a fundamental tradeoff between effectiveness (e.g., recall, MRR, nDCG) and resource efficiency (latency, throughput, memory footprint, and index/update cost). Architectural choices — how queries and documents are represented, how similarity is computed, and how candidates are staged and scored — drive where a method falls on this tradeoff curve. Modern neural retrievers demonstrated that learned dense embeddings can substantially improve effectiveness over classic lexical baselines, but achieving that gain requires extra compute and index engineering compared to lightweight lexical methods.
A common high-performance design is the bi-encoder (single-vector) architecture (Sections˜6 and 7): queries and documents are encoded independently into compact vectors and approximate nearest-neighbor search (ANN) retrieves candidates quickly. This architecture is attractive for strict latency budgets and large corpora because it enables highly optimized ANN indexes (e.g., HNSW (Malkov & Yashunin, 2018)) that deliver millisecond-scale queries at large scale; however, single-vector representations can miss fine-grained token-level matches that matter for some queries.
To improve effectiveness, researchers have pursued richer interaction patterns. Late-interaction or multi-vector models (e.g., ColBERT and ColPali (Khattab & Zaharia, 2020; Faysse et al., 2025)) keep token-granular signals and aggregate local token similarities, which raises effectiveness but increases index size and per-query compute; compression and residual quantization techniques can reduce the space/latency penalty but do not eliminate it entirely. Similarly, learned sparse/lexical models (e.g., SPLADE (Formal et al., 2021a; b)) produce high-dimensional but sparse representations that recover lexical signals and interpretability while trading off somewhat higher compute or indexing complexity versus classical BM25. These architectural variants illustrate the spectrum: more expressive interactions often translates to better quality, typically at higher memory and latency cost (unless mitigated by compression).
A pragmatic pattern in production is staged (two- or multi-stage) retrieval: a fast, coarse first-stage (BM25 or compact dense ANN) produces a small candidate set, and a more expensive cross-encoder or interaction-based reranker refines the top results (Huang et al., 2020; Su et al., 2025). This cascade yields most of the accuracy of expensive models while preserving throughput, but it requires careful budgeting (how many candidates to pass) and engineering (batching, caching, and efficient GPU/CPU placement). Indexing and compression techniques (e.g., product quantization, pruning, residual quantization), along with hybrid lexical–dense pipelines, are commonly employed to navigate the tradeoff curve as operational constraints evolve.
Architectural Design Heuristics.
Adopt a single-vector dual-encoder with efficient ANN search when low latency, low cost, and frequent updates are the primary constraints. Favor multi-vector or interaction-intensive models when top-tier ranking quality is required and computational budget permits. To balance these objectives, employ a two-stage pipeline comprising coarse retrieval followed by a specialized reranker. In large-scale deployments, the effectiveness-efficiency operating point is chiefly adjusted through index compression techniques, hybrid lexical-dense signals, and careful calibration of candidate set size.
8.3 Calibration and Confidence Estimation
Calibration in IR aims to align model scores with true probabilities of relevance, such that confidence faithfully reflects correctness. Unlike traditional deterministic rankers that output single relevance scores, calibrated IR models explicitly represent uncertainty, often as a distribution over scores whose mean encodes relevance belief and whose variance captures uncertainty (Cohen et al., 2021). This distinction is architecturally significant: uncertainty-aware scoring exposes information that is otherwise hidden in standard ranking pipelines. The importance of calibration is formalized by the Probability Ranking Principle (PRP), which guarantees optimal ranking only when relevance probabilities are well calibrated and reported with certainty (Penha & Hauff, 2021). However, modern neural rankers frequently violate these assumptions, motivating calibration-aware architectural design.
Neural Architectures and Uncertainty Modeling.
Calibration properties in IR are strongly dependent on architectural choices, with empirical studies showing mixed calibration behavior across neural model families (Guo et al., 2017; Minderer et al., 2021). Transformer-based rankers, including BERT variants, are often poorly calibrated, with calibration quality varying by model scale and design (Dan & Roth, 2021; Li et al., 2022a). Introducing stochasticity at the architectural level — through stochastic inference or approximate Bayesian formulations — consistently improves calibration compared to deterministic counterparts, while adding only modest computational overhead (Penha & Hauff, 2021; Cohen et al., 2021). These findings position stochastic and Bayesian architectures as principled mechanisms for embedding uncertainty directly into relevance estimation.
Calibration in Multi-Component Retrieval Systems.
Calibration challenges are amplified in multi-stage retrieval architectures, such as the “retrieve-then-rerank” pipeline, where hard top-$k$ retrieval steps break differentiability and preclude end-to-end calibration. Architectural interventions, including differentiable sampling mechanisms based on Gumbel approximations, restore gradient flow and enable joint calibration of retriever and reader components (Dhuliawala et al., 2022). Jointly calibrated systems produce more reliable confidence estimates than calibrating individual modules in isolation, underscoring how calibration requirements can directly shape architectural design in complex IR pipelines.
Modularity and Calibration-Aware Computation.
Beyond end-to-end design, modular architectures enable calibration to be treated as an attachable component rather than an intrinsic model property. Universal post-hoc calibrators can be applied across heterogeneous retrieval architectures without modifying their internals, offering scalable and architecture-agnostic improvements (Zhang et al., 2021a). When uncertainty is explicitly modeled, calibration becomes operational rather than merely diagnostic: confidence estimates can guide adaptive computation, such as selective reranking or deferred inference for ambiguous queries, improving both efficiency and robustness (Cohen et al., 2021; Yoon et al., 2025). Recent evidence further suggests that architectural specialization — e.g., modeling relevance across multiple criteria instead of a single scalar score — can inherently reduce calibration error, reinforcing calibration as a first-class architectural consideration in IR system design (Penha & Hauff, 2021; Javdan et al., 2025). To summarize, how to incorporate calibration into modern retrieval systems is still an open question in the IR community.
8.4 Domain-Specific Applications
General Web Search.
General-purpose web search has historically been dominated by the “retrieve-then-rerank” paradigm, built around inverted indexes and multi-stage cascade ranking to ensure efficiency at web scale (Shen et al., 2014a; Mitra et al., 2016; Qin et al., 2022; Zhang et al., 2025a). These architectures decompose retrieval into heterogeneous components for query understanding, candidate generation, ranking, and re-ranking, each optimized independently (Wang & Na, 2023). As web content diversified, search engines incorporated additional signals such as anchor text, hyperlinks, and layout-based features to enrich document representations beyond plain text (Oliveira & Teixeira Lopes, 2023). While highly effective, these pipelines impose rigid, predefined information flows that limit adaptability. The recent generative retrieval (Section˜7.3) thus explores replacing this modular pipeline with model-centric architectures, where a single LLM performs indexing, retrieval, and ranking end-to-end.
Domain-Specific and Thematic Retrieval.
Domain-specific search engines target focused corpora such as scientific literature, medicinal chemistry databases, and job postings, leveraging domain knowledge to improve relevance beyond general web search. These applications share architectural challenges arising from specialized terminology, underspecified expert queries, and narrowly scoped user intents (Wang & Na, 2023; Kang et al., 2024). Generic pretrained models often fail to capture domain-specific semantics without architectural support for structured knowledge. Traditional solutions rely on query expansion, lexical analysis, and large task-specific training sets, but suffer from scalability and complexity limitations (Xu et al., 2025f). In response, modern architectures increasingly integrate entity- and relation-centric representations to better align retrieval with domain semantics (Dong et al., 2022a). As a result, domain-specific IR architectures emphasize knowledge-aware representations and domain adaptation over purely generic pretrained models.
Medical and Legal Information Retrieval.
Medical and legal search systems represent high-stakes instantiations of domain-specific retrieval, imposing stricter architectural constraints. Medical IR must operate over long temporal records that combine unstructured clinical notes with structured diagnoses, laboratory values, and medications, motivating architectures that integrate information extraction, faceted search, and structured database querying (Sonntag & Profitlich, 2018). These systems must additionally handle specialized medical terminology, clinical documentation standards, and stringent privacy requirements. Legal IR faces analogous challenges due to highly specialized language, complex document structures, and limited labeled relevance data (Althammer et al., 2020). Despite recent efforts to adapt pretrained language models such as BERT, neural approaches often fail to consistently outperform strong lexical baselines like BM25, reinforcing the continued relevance of hybrid and lexically grounded architectures in legally specialized domains (de Araujo et al., 2014; Althammer et al., 2020).
E-commerce Search.
E-commerce search systems operate over structured product catalogs rather than unstructured web pages, fundamentally shaping their architectures (Wang & Na, 2023; Ren et al., 2024). The dominant approach employs embedding-based retrieval using the bi-encoder architecture combined with approximate nearest neighbor search (Nigam et al., 2019; Lin et al., 2024). These systems must address severe vocabulary mismatch between customer queries and product descriptions, motivating query rewriting and semantic bridging components, often implemented using LLMs trained with contrastive or instruction-based objectives (Peng et al., 2024b).
Ads Ranking.
Ads ranking is another important industrial IR application scenario. Modern advertising recommendation systems face the fundamental challenge of processing massive candidate sets while maintaining strict latency requirements and ranking quality. The traditional approach employs multi-stage cascading architectures that decompose the Ad ranking problem into sequential stages: retrieval, pre-ranking (or lightweight ranking), ranking (or heavyweight ranking), and auction (Gallagher et al., 2019; Wang et al., 2023b; Zheng et al., 2024a; Yang et al., 2025d), where the research focus is similar to that of ad hoc retrieval.
In the context of Ads ranking, the XMC formulation described in Section˜5 has been shown to improve ranking quality of tail items/Ads (Jain et al., 2019; Dahiya et al., 2021; Yu et al., 2022a; Dahiya et al., 2023; Gupta et al., 2024b), which is often critical for revenue maximization. However, it also comes with higher engineering complexity and often requires retraining of the models when new items/Ads are added.
The recent breakthrough of LLMs has catalyzed a paradigm shift toward generative recommendation frameworks that can directly generate personalized item sequences from user interaction histories. These approaches index items with meaningful IDs using vector quantization algorithms and generate items from the entire item set for recommendation, conceptualizing online advertising systems as a unified generative process that eliminates inherent goal conflicts between different pipeline stages (Rajput et al., 2023; Zheng et al., 2023; Zhai et al., 2024). Given space constraints, we refer readers to the cited works for a more in-depth treatment of the field.
Cross-Application IR.
Across applications, several architectural trends are reshaping IR system design. Retrieval-augmented generation (RAG) architectures combine dense vector search with LLMs to support search, reasoning, and synthesis over structured and unstructured data (Asai et al., 2024c). Modular and multi-agent retrieval frameworks further decouple retrieval functions into interoperable components, enabling dynamic adaptation to application-specific requirements (Chang et al., 2025; Zhang et al., 2025a). Finally, contextual personalization and the convergence of search and recommendation architectures reflect a broader shift toward unified representations of users, documents, and intent (Zamani & Croft, 2018).
Taken together, the diversity of application-driven architectures underscores that modern IR systems are no longer optimized solely for static relevance ranking, but must increasingly adapt to new roles, constraints, and integration patterns that give rise to emerging architectural directions and open challenges.
@@ 9 Emerging Directions and Challenges
IR systems have become crucial across diverse domains, from retrieval-augmented language modeling (Khandelwal et al., 2020; Borgeaud et al., 2022) to applications in agents (Wu et al., 2023a; Wang et al., 2024a; 2025a), code generation (Wang et al., 2025b; Zhang et al., 2023a), robotics (Anwar et al., 2024), medicine (Jeong et al., 2024), and protein research (Jumper et al., 2021), inter alia. These developments present new challenges for IR research. Drawing from the evolution of IR architectures (Sections˜3, 4, 5, 6, 7 and 8), we examine emerging trends, open problems, and potential research directions. Our discussion is structured around three key areas: advancing the core components of IR models, adapting to the new paradigm of retrieval for AI, and tackling the pragmatic challenges of real-world deployment.
@@ 9.1 Advancing the Core Components of IR Models
  • Data and training efficiency. Current transformer-based IR models demand extensive training data (Fang et al., 2024), making them impractical for many real-world applications. Developing architectures that can learn effectively from limited data or in a few-shot/zero-shot setting remains crucial. Additionally, models should support parallel processing and low-precision training to reduce costs and accelerate convergence (Pool et al., 2021; Fishman et al., 2024; Liu et al., 2024a).
  • Inference optimization. Real-time applications like conversational search (Mo et al., 2024b) and agent-based systems (Yao et al., 2023) require efficient handling of variable-length queries, necessitating advanced compression and optimization techniques for both retriever backbones and index structures (Dettmers & Zettlemoyer, 2023; Kumar et al., 2024; Xu et al., 2024b; Warner et al., 2025; Bruch et al., 2024; Xu et al., 2025b, inter alia).
  • Better lite foundation models. IR models need to process queries in real time, and using compact-sized foundation models is often a practical solution. Warner et al. (2025) presented an interesting study on wide-and-shallow versus deep-and-narrow architecture in the context of training a “modern” BERT model. Recent works (Günther et al., 2024; Fu et al., 2023; Portes et al., 2023) studied training efficient BERT model that supports longer context length, while Nussbaum & Duderstadt (2025) investigated the efficacy of a mixture-of-expert BERT-style encoder model. The best recipe for lite foundational models for IR applications remains an open question.
  • Transformer alternatives. While transformers have dominated recent IR research, their quadratic complexity in attention computation remains a significant bottleneck. Recent advances in linear RNNs (Peng et al., 2023; 2024a; Qin et al., 2024c), state space models (Gu & Dao, 2024; Dao & Gu, 2024), and linear attention (Katharopoulos et al., 2020; Yang et al., 2025c) offer alternatives with theoretical linear complexity. Although preliminary studies (Xu, 2024; Zhang et al., 2024; Xu et al., 2025e; c) show limited gains compared to optimized transformers, developing efficient alternative architectures for transformers could revolutionize large-scale IR.
Ultimately, strong foundation models have proven crucial for IR performance (Neelakantan et al., 2022; Ma et al., 2024b, inter alia). As IR applications expand, developing foundation models that balance computational efficiency with robust performance across tasks and modalities emerges as a key research priority.
@@
As discussed in Section˜6, cross-encoders provide complex non-linear relevance estimation but are computationally expensive. In contrast, bi-encoder architectures used in dense and sparse retrieval rely on linear similarity functions (e.g., inner product) to enable fast retrieval through nearest neighbor search and inverted indexing. Balancing complex relevance matching and scalable retrieval remains challenging. ColBERT (Khattab & Zaharia, 2020) addresses this by using document representation matrices with MaxSim operations, while recent work (Killingback et al., 2025) explores Hypernetworks (Ha et al., 2022) to generate query-specific neural networks for relevance estimation. The design of flexible yet scalable relevance estimators remains an active research direction.